Complete the equation of the line L satisfying each of the following sets of geometric conditions by determining the value of b1 and b0.
L passes through point (0,-4) and is parallel with the line 3x - y = 6. (y = b1x + b0) I got b1=-3 and b0= -4. Is this right.

Answer by stanbon(75887)

L passes through point (0,-4) and is parallel with the line 3x - y = 6.
------------------
Put the given equation in the slope-intercept form:
3x - y = 6
y = 3x -6
slope = 3
---------------
Any line parallel to the given line has slope = 3.
---------------
For of the equation you want is y = mx + b
m = 3 and y = -4 when x = 0 ; solve for "b":
-4 = 3*0 + b
b = -4
-------------
EQUATION:
y = 3x - 4
b1 is 3
b0 is -4
==============
Cheers,
Stan H.